# Limits of Safe AI Deployment: Differentiating Oversight and Control

### v1

@@ Abstract
Oversight and control, which we collectively call supervision, are often invoked as key levers for ensuring that AI systems are accountable, reliable, and able to fulfill governance and management requirements. However, the concepts are frequently conflated or insufficiently distinguished in academic and policy discourse, undermining efforts to design or evaluate systems that should remain under meaningful human supervision.
@@
This paper undertakes a targeted critical review of literature on supervision outside of AI, along with a brief summary of past work on the topic related to AI. We then differentiate control as being ex-ante or real-time, and operational rather than policy or governance. In contrast, oversight is either a policy and governance function, or is ex-post. We suggest that control aims to prevent failures. In contrast, oversight often focuses on detection, remediation, or incentives for future prevention; all preventative oversight strategies nonetheless necessitate control.
@@
Building on this foundation, we make three contributions. First, we propose a theoretically-informed yet policy-grounded framework that articulates the conditions under which each mechanism is possible, where they fall short, and what is required to make them meaningful in practice. Second, we outline how supervision methods should be documented and integrated into risk management, and drawing on the Microsoft Responsible AI Maturity Model, we outline a maturity model for AI supervision. Third, we explicitly highlight some boundaries of these mechanisms, including where they apply, where they fail, and where it is clear that no existing methods suffice. This foregrounds the question of whether meaningful supervision is possible in a given deployment context, and can support regulators, auditors, and practitioners in identifying both present limitations and the need for new conceptual and technical advances.
Background and Literature Review
Oversight and control are frequently invoked as mechanisms for ensuring AI system safety, yet their meaning is often unclear, inconsistent, or disputed. As an initial step, we reviewed both management literature and AI-specific discussions of the terms before attempting to provide clarity.
@@
In the AI context, Tsamados, Floridi, and Taddeo (2024) discusses supervisory control versus teaming as forms of “operational human control” during deployment, using Verdiesen, Santoni de Sio, and Dignum (2021)’s framework for autonomous weapons systems, as shown in Figure 3. This differentiates timing of control (before, during, or after deployment) and governance. The time dimension has an obvious meaning for weapons systems, where the deployment is a fixed, presumably short time window, but the corresponding time frames for any AI system can be understood similarly. That is, before involves design decisions, during involves operational decisions, and after involves review. In the last case, even if the system is still running, post-deployment would involve looking at the system’s past actions, rather than control and changing the outputs. However, Tsamados, Floridi, and Taddeo (2024) focuses on timing to the exclusion of Verdiesen, Santoni de Sio, and Dignum (2021)’s critical dimension distinguishing governance, socio-technical, and engineering layers of control—central for differentiating methods of oversight and control.
@@
Finally, Bengio (2024), in his US Senate testimony, focused on catastrophic outcomes and underscored four areas where human choices shape AI risk; ”(1) access—who can tinker with powerful AIs, what protocols must they follow, under what kind of oversight? (2) misalignment—the extent to which powerful AIs act not as intended by their developers, potentially causing severe or even catastrophic harm;” as well as “(3) raw intellectual power… and (4) scope of actions—the ability to affect the world and cause harm in spite of society’s defenses.“ However, once again the technical question of how oversight is accomplished and whether it is effective is left unanswered.
With that background, and keeping in mind the problem of proposing new idea in already saturated fields (Munroe 2011), we introduce domain-specific definitions of each of the two terms. We group them together as supervision - not to suggest that this additional term is useful on its own, but as a collective term for both methods for the below discussion.
@@
Using the above frameworks and varied definitions, we propose that control can only refer to actions taken before or during deployment, and to management and technical methods, as show in Figure 3 in red.
@@
The simple cases are those discussed in the diagram directly, and we group technical and socio-technical as operational mechanisms, rather than policy or governance mechanisms. Similarly, we exclude governance-level mechanisms from control. This means that control follows the contours of how the term is used in control theory. (Ogata 2010) This partly includes work on what has recently been styled ”AI Control” (Greenblatt et al. 2024) but our definition includes more prosaic risks of control failure and simpler systems, and excludes red-teaming and similar approaches proposed there, which we classify as oversight.
Note that there are also ways to build systems in which the output is controlled by some external constraint which function as control, and this could be interpreted as an ex-post control measure - though as an operational function, it cannot be a governance intervention. However, we will suggest below that when such an external constraint is functioning as a control, it is better understood as an ex-ante strategy implemented at either the technical or socio-technical level for the broader system that includes that external constraint. Similarly, we argue that what control theorists would call post-deployment feedback mechanisms are either put in place prior to deployment, or are later modifications triggered by a governance or oversight decision.
@@
Time-scope and Type -
Ex-ante, In-flight, or Ex-post, and type of control (e.g., “Control - Technical”).
Purpose / Risks Mitigated -
@@
Table 1: Control Mechanisms: Summary + Details
Method Time-Scope Possible Risks Mitigated Notes
Technical Control Ex-ante Illegal outputs; specified narrow harms 1
Human Control Ex-ante, In-flight Automated mistakes; shutdown 2
1
Examples: Programmed constraints, sandboxes, rule‑based limits, control systems, formal verification
Enablers: fully specified environment & provable world‑model, sufficient observability.
Failure Modes: Specification gaps; control system failures; adversarial bypass.
2
Examples: Human-in-the-loop; human control via variable‑autonomy protocols; shutdown and lock‑out procedures for worst-case risks, e.g. Mickens, Radway, and Netravali (2025).
Enablers: Human attention, aligned incentives, training, interpretable and/or explainable AI systems, and adequate human‑machine supervision interfaces, ability to identify when loss-of-control occurs.
Failure Modes: Alert fatigue and failures of attention, expertise, and/or situational awareness; skill atrophy, routine-ization, and boredom; inability to detect failures rapidly enough
Table 2: Oversight Mechanisms: Summary + Details
Method Type Time-Scope Possible Risks Mitigated Notes
Preventive Active Ex-ante Uncertain dangerous capabilities; identify exploits; ensuring design meets specifications. 1
Human-above-the-Loop In-flight Quickly-rectifiable mistakes; flagged misbehavior. 2
Management Review Ex-post Rectifiable failures; bias; non-immediate failure modes. 3
Internal Governance Ex-post Long-term failures; identification of previously unidentified risks. 4
1
Examples: Red-teaming; predeployment audits; internal and external review.
Enablers: Management and cross-stakeholder buy-in; technical expertise; sufficient time before deployment; precise risk definitions.
Failure Modes: Failure to anticipate failures; treadmill effect as models improve and diagnosis methods fall behind; excessive focus on easily tested failure modes.
2
Examples: Tiered escalation; adaptive autonomy.
Enablers: Dynamic policy; continuous monitoring; clear definition of failures.
Failure Modes: Institutional inertia; slow response time; inability to identify or rectify failures.
3
Examples: Logging and anomaly detection; postmortems; management reporting; public disclosures.
Necessary enablers: Reliable instrumentation; resources for assurance and audit; ombudsman and whistle-blower protection; strong safety culture; independent reporting channels and escalation pathways.
Failure Modes: Detection latency; irreversibility; internal pressure; lack of transparency; liability concerns; incorrect root-cause diagnosis.
4
Examples: Review and audit systems; access restrictions.
Enablers: Robust safety culture; external accountability; enforceability.
Failure Modes: Jurisdictional gaps; lack of enforcement; routinization.
5
Examples: Liability regimes; government/international oversight; third-party compliance reviews.
Enablers: Established in advance; early external knowledge of risks and failures.
Failure Modes: Detection latency; irreversibility.
@@
Second, metrics are important for management in general, and oversight in particular, but developing useful such metrics pose particular challenges (Manheim 2023b). The development of metrics for GenAI models themselves is still a rapidly evolving field (Weidinger et al. 2025), and the development of metrics for meaningful oversight is in many ways even more rudimentary. This is critically different from measurement of the risks themselves, but likely requires progress on that front as well. That said, even simple metrics, like Coverage $\times$× Detection $\times$× Remediation rates, can be useful in situations where no other quantitative process for measuring or evaluation of supervision exists.
@@
For control and some oversight methods, we can have issues with speed mismatches, when time‑to‑harm is less than the control system reaction time, or in the case of human involvement, less than the time needed to understand and react. There are also problems with under-specified or ambiguous contexts that can include novel scenarios outside training or procedure scope. For oversight, failure modes range from limited human capacity, including attention, expertise, situational awareness, to organizational incentives such as against raising red flags. Notably, the organizational failures for some oversight methods need not be within the organization, such as reliance on oversight bodies influenced by industry or relying on external laws which do not keep pace with capabilities. Similarly, technical oversight issues can include a failure of technical transparency, such as insufficient interpretability and missing or limited audit methods or tools, insufficient logging or log analysis.
@@
A supervision process that outlines the mitigations and risks addressed clearly, and implements an effective oversight process is ideal. Of course, it is not always the case that the ideal is realized, and while successful prevention of a harm is binary, strategies and management approaches to mitigate the risks are not. For that reason, we suggest a maturity model for AI supervision. This should be useful, for example, for system designers making claims about their mitigations, and for auditors or regulators assessing the systems.
In Table 3, we suggest an adaptation of one dimension of the Microsoft Responsible AI Maturity Model, that for mitigating AI risks. (Vorvoreanu et al. 2023)
@@
5 Teams implement supervisory or control mitigations for all appropriate risks identified internally, as well as those disclosed in similar systems. Along with clear publicly available risk identification, there is public documentation of mitigations before even internal deployment, including monitoring and risk prioritization. This also acknowledges unknowns and monitors for new uses, risks, and system changes, and prepares to adapt to residual risk. completeness
@@ Conclusion
Not all risks from AI can be resolved with oversight or control. Similarly, many ethical dilemmas posed by AI are not be resolvable through better oversight mechanisms alone and might require societal discussions about whether certain AI applications should be developed or deployed at all. Despite this, oversight and control play a useful role, and hopefully the above discussion enables better discussion, design, and operational management of supervision systems. The requirements for different strategies highlight the difficulty, especially if used in real-time and at scale, and when the system itself is poorly understood. It should therefore be clear that meaningfully controlling or supervising a given system might not be possible.
For this reason, our conclusion is hopeful for the better management of limited and well-understood systems, and to address the classes of risks which can be addressed. Outside of those systems, we reiterate that we advise caution against the idea that either control or oversight is a viable risk-mitigation strategy for many classes of proposed future autonomous systems which pose unknown-unknown risks, much less a sufficient solution for ensuring their safety.
Note: LLMs were used for feedback on the ideas before, during, and after the writing, and for copy-editing for grammar and similar issues. They were also used extensively for LaTeX formatting, suggesting table and figure formats, and figure generation, and generating BibTeX citations, which were reviewed.

### v2

@@ Abstract
Oversight and control, which we collectively call supervision, are often discussed as ways to ensure that AI systems are accountable, reliable, and able to fulfill governance and management requirements. However, the requirements for ”human oversight” risk codifying vague or inconsistent interpretations of key concepts like oversight and control. This ambiguous terminology could undermine efforts to design or evaluate systems that must operate under meaningful human supervision. This matters because the term is used by regulatory texts, such as the EU AI Act, the NIST AI Risk Management Framework, and the OECD AI Principles.
@@
This paper undertakes a targeted critical review of literature on supervision outside of AI, along with a brief summary of past work on the topic related to AI. We then differentiate control as being ex-ante or real-time, and operational rather than policy or governance. In contrast, we view oversight as either performed ex-post, or a policy and governance function. We suggest that control aims to prevent failures, while oversight focuses on detection, remediation, or incentives for future prevention, and we note that preventative oversight strategies necessitate control.
@@
Building on this foundation, we make three contributions. First, we propose a framework to align regulatory expectations with what is technically and organizationally plausible, articulating the conditions under which each mechanism is possible, where they fall short, and what is required to make them meaningful in practice. Second, we outline how supervision methods should be documented and integrated into risk management, and drawing on the Microsoft Responsible AI Maturity Model, we outline a maturity model for AI supervision. Third, we explicitly highlight some boundaries of these mechanisms, including where they apply, where they fail, and where it is clear that no existing methods suffice. This foregrounds the question of whether meaningful supervision is possible in a given deployment context, and can support regulators, auditors, and practitioners in identifying both present and future limitations.
Background
Oversight and control are frequently invoked as mechanisms for ensuring AI system safety, including in regulation, yet the terms are conflated or used interchangeably, in part because their meaning is often unclear, inconsistent, or disputed. Clarifying this is critical for legal requirements and standards which require either, or both. While this is a seemingly narrow component of governance, reliance on oversight and control as a central part of governance is nearly universal, including in American policy (Executive Office of the President 2023) and risk management frameworks (National Institute of Standards and Technology (2023) NIST), in the EU AI Act, (European Parliament and Council of the European Union 2024), in OECD principles for AI, (OECD 2024) and in various international standards. (iso 2022, 2023)
We begin by considering the terms and what they mean based on a review of both management literature and AI-specific discussions of the terms. The review leads to a conclusion that control refers to operational management and technical methods taken before or during deployment to prevent failures. In contrast, oversight can be passive, post-hoc operational processes or governance methods. It can also be active, that is, methods which enable human control to prevent failure.
This clarity about meaning allows us to provide further explanation the processes which are used to accomplish the goals of supervision of AI systems, and how they must be reported. Finally, we provide a maturity model to indicate the extent to which the goals of supervision of AI systems are accomplished. Hopefully, these three pieces - clarity about what is required, reporting of what is done, and evaluation of success - provide the necessary conditions needed for clear standards, effective regulation, and eventually, safe AI systems.
Literature Review
To start, we note that supervision is a governance requirement that applies in many domains, and is only one of several necessary approaches, as illustrated in figure 1. Nonetheless, AI Governance has specific issues, so our review seeks to understand the ideas first in the context of management and governance generally, then within the specific context of AI system supervision.
@@
In the AI context, Tsamados, Floridi, and Taddeo (2024) discusses supervisory control versus teaming as forms of “operational human control” during deployment, using Verdiesen, Santoni de Sio, and Dignum (2021)’s framework for autonomous weapons systems, as shown in Figure 3. This differentiates timing of control (before, during, or after deployment) and governance. The time dimension has an obvious meaning for weapons systems, where the deployment is a fixed, presumably short time window, but the corresponding time frames for any AI system can be understood similarly. That is, before involves design decisions, during involves operational decisions, and after involves review. In the last case, even if the system is still running, post-deployment would involve looking at the system’s past actions, rather than control and changing the outputs. However, Tsamados, Floridi, and Taddeo (2024) focuses on timing to the exclusion of Verdiesen, Santoni de Sio, and Dignum (2021)’s critical dimension distinguishing governance, sociotechnical, and engineering layers of control—central for differentiating methods of oversight and control.
@@
Finally, Bengio (2024), in his US Senate testimony, focused on catastrophic outcomes and underscored four areas where human choices shape AI risk; ”(1) access—who can tinker with powerful AIs, what protocols must they follow, under what kind of oversight? (2) misalignment—the extent to which powerful AIs act not as intended by their developers, potentially causing severe or even catastrophic harm” as well as “(3) raw intellectual power… and (4) scope of actions—the ability to affect the world and cause harm in spite of society’s defenses.“ However, once again the technical question of how oversight is accomplished and whether it is effective is left unanswered.
With that background, and keeping in mind the problem of proposing new ideas in already saturated fields (Munroe 2011), we introduce domain-specific definitions of each of the two terms, and provide a typology which is outlined in Figure 2. We group them together as supervision - not to suggest that this additional term is useful on its own, but as a collective term for both methods for the below discussion.
@@
Using the above frameworks and varied definitions, control can only refer to actions taken before or during deployment, and to management and technical methods, as shown in Figure 3 in red.
@@
The simple cases are those discussed in the diagram directly, and we group technical and sociotechnical as operational mechanisms, rather than policy or governance mechanisms. Similarly, we exclude governance-level mechanisms from control. This means that control follows the contours of how the term is used in control theory. (Ogata 2010) This partly includes work on what has recently been styled ”AI Control” (Greenblatt et al. 2024) but our definition includes more prosaic risks of control failure and simpler systems, and excludes red-teaming and similar approaches proposed there, which we classify as oversight.
Note that there are also ways to build systems in which the output is controlled by some external constraint which function as control. Such external control can be an ex-post control measure - though as an operational function, it cannot be a governance intervention. However, it is better understood within the broader system that includes that external constraint as an ex-ante strategy implemented at either the technical or sociotechnical level. Similarly, we argue that what control theorists would call post-deployment feedback mechanisms are either put in place prior to deployment, or are later modifications triggered by a governance or oversight decision.
@@
Time-scope and Type
Ex-ante, In-flight, or Ex-post, and type of control (e.g., “Control – Technical”).
Purpose / Risks Mitigated
@@
Table 1: Control Mechanisms
Method Time-Scope Examples Enablers Failure Modes
Technical Control Ex-ante Programmed constraints; sandboxes; rule‑based limits; control systems; formal verification Fully specified environment & provable world‑model; sufficient observability Specification gaps; control system failures; adversarial bypass
Human Control Ex-ante, In-flight Human-in-the-loop; variable‑autonomy protocols; shutdown/lock‑out procedures for worst-case risks (e.g., Mickens, Radway, and Netravali (2025)) Human attention; aligned incentives; training; interpretable/explainable AI; adequate human‑machine supervision interfaces; ability to identify loss‑of‑control Alert fatigue; failures of attention/expertise/situational awareness; skill atrophy & routinization; boredom; slow failure detection
Table 2: Oversight Mechanisms
Method Type Time-Scope Examples Enablers Failure Modes
Preventive Active Ex-ante Red-teaming; predeployment audits; internal/external review Management & cross-stakeholder buy-in; technical expertise; time before deployment; precise risk definitions Failure to anticipate failures; treadmill effect as models improve; focus on easily tested failure modes
Human-above-the-Loop In-flight Tiered escalation; adaptive autonomy Dynamic policy; continuous monitoring; clear failure definitions Institutional inertia; slow response; inability to identify or rectify failures
Management Review Ex-post Logging & anomaly detection; postmortems; management reporting; public disclosures Reliable instrumentation; assurance/audit resources; ombuds/whistle-blower protection; strong safety culture; independent reporting & escalation channels Detection latency; irreversibility; internal pressure; lack of transparency; liability concerns; incorrect root-cause diagnosis
Internal Governance Ex-post Review/audit systems; access restrictions Robust safety culture; external accountability; enforceability Jurisdictional gaps; lack of enforcement; routinization
External / Regulatory Oversight Ex-post Liability regimes; government/international oversight; third-party compliance reviews Established in advance; early external knowledge of risks & failures Detection latency; irreversibility
Second, metrics are important for management in general, and oversight in particular, but developing useful such metrics pose particular challenges (Manheim 2023b). The development of metrics for GenAI models themselves is still a rapidly evolving field (Weidinger et al. 2025), and the development of metrics for meaningful oversight is in many ways even more rudimentary. This is critically different from measurement of the risks themselves, but likely requires progress on that front as well. That said, even simple metrics, like Coverage $\times$ Detection $\times$ Remediation rates, can be useful in situations where no other quantitative process for measuring or evaluation of supervision exists.
@@
For control and some oversight methods, we can have issues with speed mismatches, when time‑to‑harm is less than the control system reaction time, or in the case of human involvement, less than the time needed to understand and react. There are also problems with under-specified or ambiguous contexts that can include novel scenarios outside training or procedure scope. For oversight, failure modes range from limited human capacity, including attention, expertise, situational awareness, to organizational incentives such as against raising red flags.
Notably, the organizational failures for some oversight methods need not be within the organization, such as reliance on oversight bodies influenced by industry or relying on external laws which do not keep pace with capabilities. Similarly, technical oversight issues can include a failure of technical transparency, such as insufficient interpretability and missing or limited audit methods or tools, insufficient logging or log analysis.
@@
In practice, the idealized process we present, which outlines the mitigations and risks addressed clearly and then implements an effective oversight process, may not be achieved. While successful prevention of a harm is binary, strategies and management approaches to mitigate the risks are not. For that reason, we suggest a maturity model for AI supervision. This should be useful, for example, for system designers making claims about their mitigations, and for auditors or regulators assessing the systems.
@@
5 Teams implement supervisory or control mitigations for all appropriate risks identified internally, as well as those disclosed in similar systems. Along with clear publicly available risk identification, there is public documentation of mitigations before even internal deployment, including monitoring and risk prioritization. This also acknowledges unknowns and monitors for new uses, risks, and system changes, and prepares to adapt to residual risk.
@@
In Table 3, we suggest an adaptation of one dimension of the Microsoft Responsible AI Maturity Model, that for mitigating AI risks. (Vorvoreanu et al. 2023) This can be used alongside a broader review of overall maturity, but can also be used as an independent evaluative tool.
@@ Conclusion
In addition to clarifying terminology in ways that support standards processes and regulation, we have provided a clear path forward for better documentation and understanding. This is important for those developing, deploying, testing, or evaluating the purchase of models. However, we caution that not all risks from AI can be resolved with oversight or control. In addition, ethical dilemmas posed by AI are not resolvable through better oversight mechanisms alone. They also cannot address whether certain AI applications should be developed or deployed at all, which requires broader risk management, and societal discussions.
The requirements for different strategies highlight that such oversight is not easy, especially if used in real-time and at scale, and when the system itself is poorly understood. Therefore, meaningfully controlling or supervising a given system might not be possible. Still, oversight and control play a useful role, and the above discussion enables better discussion, design, and operational management of supervision systems.
For this reason, our conclusion is hopeful for the better management of limited and well-understood systems, and to address the classes of risks which can be addressed. Outside of those systems, we reiterate that we advise caution against the idea that either control or oversight is a sufficient solution for ensuring their safety, or even a viable risk-mitigation strategy for many proposed future autonomous systems, especially those which pose unknown-unknown risks.